OpenStudy (ellamoyseyuk):

Jack and Maureen are opening an ice hotel. The structure and furniture are made out of ice. They must determine the weight of some of the ice furniture to decide the size of the crane needed to lift each piece into place. Maureen knows that ice weighs approximately 62.4 pounds per cubic foot. Find the approximate weight of each of the figures below. a. The stage in the banquet hall measures 22 ft by 15 ft by 4 ft. b. Two cylindrical pillars will flank the front door. Each cylinder has diameter 1 ft and height 9 ft. c. A series of four spheres will decorate the reception area. The largest

OpenStudy (ellamoyseyuk):

A) V = l*w*h V = 22*15*4 V = 1320 = 1320*62.4 = 82368 pounds. B) V = Bh V = pi ½^2*9 V = pi*0.25*9 V = 2.25 pi V = 7.07 = 7.07 * 62.4 = 441.168 pounds. V = Bh V = pi ½^2*9 V = pi*0.25*9 V = 2.25 pi V = 7.07 = 7.07 *62.4 = 441.168 pounds C) v= 4/3pir^3 V = 4/3pi*4^3 v =4/3 pi*64 v = 85.33 pi v = 268.07 = 268.07 * 62.4 = 16727.568 pounds v= 4/3pir^3 V = 4/3pi*2^3 v =4/3 pi*8 v = 10.67 pi v = 33.52 = 33.52 * 62.4 = 2091.648 pounds. v= 4/3pir^3 V = 4/3pi*1^3 v =4/3 pi*1 v = 1.33 pi v = 4.18 = 4.18 * 62.4 = 260.832 pounds. v= 4/3pir^3 V = 4/3pi*1/2^3 v =4/3 pi*0.125 v = 0.17 pi v = 0.53 = 0.53 * 62.4 = 33.072 pounds.

OpenStudy (ellamoyseyuk):

is that the correct answer?

OpenStudy (foxesrcool):

There is no correct answer. The ice will melt before they are done building.

OpenStudy (ellamoyseyuk):

@pooja195

OpenStudy (ellamoyseyuk):

please can u help me

OpenStudy (ellamoyseyuk):

@welshfella